[Brahms-l] Shift report 20050313 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Mar 13 2005 - 00:06:40 EST
Shifter leader was SB
Shifters were 


1. Beam has been good the whole shift. 
2. Runs Accomplished
FS 3 deg 1723B, MRS 90 deg 700A
-------------------------------
Run     tr2   tr3
-------------------------------
13731   58k   25k
13744   8k    3k
13753   60k   25k
13754   16k   25k
13755   303k  151k
3. T1 and T2 had to be reset several times. The problem seems
to have been associated with the tap setting on D2 which was not
at Tap3. The magnet people came and fixed that and there has been
 no problem since then.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Mar 13 00:06:59 2005

This archive was generated by hypermail 2.1.8 : Sun Mar 13 2005 - 00:07:13 EST